OpenStudy (anonymous):

find gcf of 24,48 and 100

OpenStudy (anonymous):

im thinking 4...

myininaya (myininaya):

24=2*2*2*3 48=2*2*2*2*3 100=2*2*5*5 gcf(24,48,100)=2*2=4

myininaya (myininaya):

you thought right!

OpenStudy (anonymous):

^thats a good way to do it, i totally forgot about that method! :O

OpenStudy (anonymous):

what about 125 and 136

myininaya (myininaya):

write prime factorizations and see how many pairs they have in common

OpenStudy (anonymous):

i know how but im to laz.y